Use the 5-step method and be sure to include units where appropriate. Round your answers to the correct number of significant digits.
Adjust the density formula and solve the problems below:

Pure gold has a density of 119.32 g/cm3. How large would a piece of gold be if it had a mass of 3118.97 g?
1
Expert's answer
2013-09-23T09:04:48-0400
The density formula is
ρ(g/cm^3) = (m(g))/(Vcm^3))
According to this formula the volume of the metal is
V(cm^3) = □((m(g))/(ρ(g/cm^3))
V(cm^3) = 3118.97/119.32 = 26.1 cm^3

Answer: V = 26.140 cm^3